OpenStudy (anonymous):

given cos(theta)=-2/3 and theta has its terminal side in Quadrant 3, find sin(theta)

OpenStudy (jdoe0001):

\(\bf cos(\theta) = -\cfrac{2}{3} \implies \cfrac{\textit{adjacent}}{\textit{hypotenuse}} \cfrac{a}{c}\\ \textit{keep in mind that } c^2 = a^2+b^2 \implies c = \sqrt{a^2+b^2}\)

OpenStudy (jdoe0001):

so, what do you think is \(\bf sin(\theta) \ \ ?\)

OpenStudy (jdoe0001):

well, I should have said \(\bf cos(\theta) = -\cfrac{2}{3} \implies \cfrac{\textit{adjacent}}{\textit{hypotenuse}}\implies \cfrac{a}{c}\\ \textit{keep in mind that } c^2 = a^2+b^2 \implies b = \sqrt{c^2+a^2}\)

OpenStudy (jdoe0001):

acck

OpenStudy (jdoe0001):

\(\bf \textit{keep in mind that} c^2 = a^2-b^2 \implies b = \sqrt{c^2+a^2}\)

OpenStudy (jdoe0001):

sorry for the typos

OpenStudy (jdoe0001):

lemme correct myself

OpenStudy (jdoe0001):

\(\bf cos(\theta) = -\cfrac{2}{3} \implies \cfrac{\textit{adjacent}}{\textit{hypotenuse}}\implies \cfrac{a}{c}\\ \textit{keep in mind that} c^2 = a^2+b^2 \implies b = \sqrt{c^2-a^2}\)

OpenStudy (jdoe0001):

geez, much better, anyhow so what do you think will be \(\bf sin(\theta)\ \ \ ?\)

OpenStudy (anonymous):

i have no clue

OpenStudy (jdoe0001):

well, what do you think is the component "a" and "c"?